# Pivoting Factorization: A Compact Meta Low-Rank Representation of Sparsity for Efficient Inference in Large Language Models

### v1

@@ Abstract
The rapid growth of Large Language Models has driven demand for effective model compression techniques to reduce memory and computation costs. Low-rank pruning has gained attention for its tensor coherence and GPU compatibility across all densities. However, low-rank pruning has struggled to match the performance of semi-structured pruning, often doubling perplexity (PPL) at similar densities. In this paper, we propose Pivoting Factorization (PIFA), a novel lossless meta low-rank representation that unsupervisedly learns a compact form of any low-rank representation, effectively eliminating redundant information. PIFA identifies pivot rows (linearly independent rows) and expresses non-pivot rows as linear combinations, achieving an additional 24.2% memory savings and 24.6% faster inference over low-rank layers at $r/d=0.5$, thereby significantly enhancing performance at the same density. To mitigate the performance degradation caused by low-rank pruning, we introduce a novel, retraining-free low-rank reconstruction method that minimizes error accumulation (M). MPIFA, combining M and PIFA into an end-to-end framework, significantly outperforms existing low-rank pruning methods and, for the first time, achieves performance comparable to semi-structured pruning, while surpassing it in GPU efficiency and compatibility.
@@ 1 Introduction
In contrast, low-rank pruning methods, primarily based on Singular Value Decomposition (SVD), preserve the coherence of tensor shapes, making them universally compatible with any GPU architecture at any density. Recent works (Hsu et al., 2022; Yuan et al., 2023; Wang et al., 2024) have demonstrated the potential of low-rank decomposition in compressing LLMs. However, despite their flexibility, these methods struggle to compete with semi-structured pruning in terms of performance, often resulting in a 2x increase in perplexity (PPL) at the same densities. This performance gap stems primarily from two challenges: (1) SVD-based low-rank pruning introduces information redundancy in the decomposed weight matrices, and (2) existing reconstruction methods accumulate errors across layers, leading to suboptimal performance.
@@
Combining M and PIFA, we present MPIFA—an end-to-end, retraining-free low-rank pruning framework for LLMs. MPIFA significantly outperforms existing low-rank pruning methods, reducing the perplexity gap by 40%-70% on LLaMA2 models (7B, 13B, 70B) and the LLaMA3-8B model  [^1] https://www.llama.com/. Further experiments show that MPIFA consistently achieves superior speedup and memory savings both layerwise and end-to-end compared to semi-structured pruning, while maintaining comparable or even better perplexity. Notably, as shown in Table 3, at $d=32768$, PIFA with 55% density achieves a 2.1$\times$ speedup, whereas various implementations of semi-sparse methods are either slower than the dense linear layer or fail to execute. Our code is available at https://anonymous.4open.science/r/PIFA-68C3.
@@ 2.1 Connection-wise pruning
    Structured pruning (Ma et al., 2023; van der Ouderaa et al., 2024; Ashkboos et al., 2024; Lin et al., 2024) removes entire components of the model, such as neurons, channels, or attention heads, rather than individual weights. This method preserves tensor alignment and coherence, ensuring compatibility with all GPUs and enabling significant acceleration on both CPUs and GPUs. However, in LLMs, structured pruning can lead to greater loss compared to unstructured or semi-structured pruning.
@@ 2.2 Low-rank pruning
Low-rank pruning applies matrix decomposition techniques, such as Singular Value Decomposition (SVD), to approximate weight matrices with lower-rank representations, thereby reducing both storage and computational demands. This approach, compatible with any GPU, represents large matrices as products of smaller ones, improving computational efficiency. Recent studies (Hsu et al., 2022; Yuan et al., 2023; Wang et al., 2024; jai, 2024) highlight the effectiveness of low-rank decomposition in compressing LLMs. However, despite their flexibility, these methods lag behind semi-structured pruning in performance, often leading to a 2$\times$ increase in perplexity (PPL) at the same densities.
@@ 4 Online Error-Accumulation-Minimization Reconstruction
In addition to PIFA, we propose a novel Online Error-Accumulation-Minimization Reconstruction (M) method (illustrated in Figure 2) to reconstruct the low-rank matrix before applying PIFA.
@@ 5 Experiments
We combine Online Error-Accumulation-Minimization Reconstruction with Pivoting Factorization into an end-to-end low-rank compression method, MPIFA (illustrated in Figure 2). MPIFA proceeds as follows: ① First, Online Error-Accumulation-Minimization Reconstruction iis applied to obtain and refine the low-rank matrices $\vec{U}_{r}$ and $\vec{V}_{r}^{\mathrm{T}}$; ② Then, PIFA decomposes the singular matrix $\mathbf{W}^{\prime}=\vec{U}_{r}\vec{V}_{r}^{\mathrm{T}}$ into $\mathcal{I},\mathbf{W}_{p},\mathbf{C}\leftarrow\mathrm{PIFA}(\mathbf{W}^{\prime})$, which are stored in a PIFA layer that replaces the original linear layer.
@@
Compared to semi-structured sparsity, MPIFA offers the advantage of allowing non-uniform sparsity for each module. We term the Non-uniform Sparsity version of MPIFA as MPIFANS. The detail implementation of MPIFANS can be found in Appendix B.2.
@@
We apply MPIFA to pre-trained LLMs: LLaMA2 (7B, 13B, 70B) (Touvron et al., 2023b) and LLaMA3 (8B) (Dubey et al., 2024). Both the calibration data and perplexity (PPL) evaluation are based on the WikiText2 dataset (Merity et al., 2022).
@@ 5.1 Main Result
We evaluate MPIFA against state-of-the-art low-rank pruning methods: ASVD (Yuan et al., 2023) and SVD-LLM (Wang et al., 2024). Vanilla SVD is also included for reference. SVD-LLM has two versions, as detailed in their original paper. Following their approach, we select the best-performing version for each density. Results for both versions of SVD-LLM are provided in Table Pivoting Factorization: A Compact Meta Low-Rank Representation of Sparsity for Efficient Inference in Large Language Models. MPIFA utilizes 128 calibration samples and sets $\lambda=0.25$. For all models except LLaMA-2-70B, MPIFA reconstructs both $\vec{U}$ and $\vec{V}^{\mathrm{T}}$. For LLaMA-2-70B, MPIFA reconstructs only $\vec{U}$.
@@ 5.3 Ablation Study
Table Pivoting Factorization: A Compact Meta Low-Rank Representation of Sparsity for Efficient Inference in Large Language Models presents an ablation study that evaluates the impact of our Online Error-Accumulation-Minimization Reconstruction (denoted as M) and Pivoting Factorization (PIFA) on perplexity across varying parameter densities. The methods compared include:
@@
Figures 7a and 7b compare the effects of reconstructing only $\mathbf{U}$, only $\vec{V}^{\mathrm{T}}$, and both $\mathbf{U}$ and $\vec{V}^{\mathrm{T}}$ across different calibration sizes. The results indicate that with sufficient calibration samples, reconstructing both $\mathbf{U}$ and $\vec{V}^{\mathrm{T}}$ achieves lower perplexity than reconstructing only $\mathbf{U}$ or only $\vec{V}^{\mathrm{T}}$.

### v2

@@ Abstract
The rapid growth of Large Language Models has driven demand for effective model compression techniques to reduce memory and computation costs. Low-rank pruning has gained attention for its GPU compatibility across all densities. However, low-rank pruning struggles to match the performance of semi-structured pruning, often doubling perplexity at similar densities. In this paper, we propose Pivoting Factorization (PIFA), a novel lossless meta low-rank representation that unsupervisedly learns a compact form of any low-rank representation, effectively eliminating redundant information. PIFA identifies pivot rows (linearly independent rows) and expresses non-pivot rows as linear combinations, achieving 24.2% additional memory savings and 24.6% faster inference over low-rank layers at rank = 50% of dimension. To mitigate the performance degradation caused by low-rank pruning, we introduce a novel, retraining-free reconstruction method that minimizes error accumulation (M). MPIFA, combining M and PIFA into an end-to-end framework, significantly outperforms existing low-rank pruning methods, and achieves performance comparable to semi-structured pruning, while surpassing it in GPU efficiency and compatibility. Our code is available at https://github.com/biomedical-cybernetics/pivoting-factorization.
@@ 1 Introduction
In contrast, low-rank pruning methods, primarily based on Singular Value Decomposition (SVD), preserve the coherence of tensor shapes, making them universally compatible with any GPU architecture at any density. Recent works (Yuan et al., 2023; Wang et al., 2024) have demonstrated the potential of low-rank decomposition in compressing LLMs. However, despite their flexibility, these methods struggle to compete with semi-structured pruning in terms of performance, often resulting in a 2x increase in perplexity (PPL) at the same densities. This performance gap stems primarily from two challenges: (1) Low-rank pruning introduces information redundancy in the decomposed weight matrices, and (2) existing reconstruction methods accumulate errors across layers, leading to suboptimal performance.
@@
Combining M and PIFA, we present MPIFA—an end-to-end, retraining-free low-rank pruning framework for LLMs. MPIFA significantly outperforms existing low-rank pruning methods, reducing the perplexity gap by 40%-70% on LLaMA2 models (7B, 13B, 70B) and the LLaMA3-8B model  [^1] https://www.llama.com/. Further experiments show that MPIFA consistently achieves superior speedup and memory savings both layerwise and end-to-end compared to semi-structured pruning, while maintaining comparable or even better perplexity. Notably, as shown in Table 3, at $d=32768$, PIFA with 55% density achieves a 2.1$\times$ speedup, whereas various implementations of semi-sparse methods are either slower than the dense linear layer or fail to execute.
@@ 2.1 Connection-wise pruning
    Structured pruning (Ma et al., 2023; van der Ouderaa et al., 2024; Ashkboos et al., 2024; Lin et al., 2024; Song et al., 2024; Men et al., 2024; Gao et al., 2024) removes entire components of the model, such as neurons, channels, or attention heads, rather than individual weights. This method preserves tensor alignment and coherence, ensuring compatibility with all GPUs and enabling significant acceleration on both CPUs and GPUs. However, in LLMs, structured pruning can lead to greater loss compared to unstructured or semi-structured pruning.
@@ 2.2 Low-rank pruning
Low-rank pruning applies matrix decomposition techniques, such as Singular Value Decomposition (SVD), to approximate weight matrices with lower-rank representations, thereby reducing both storage and computational demands. This approach, compatible with any GPU, represents large matrices as products of smaller ones, improving computational efficiency. Recent studies (Hsu et al., 2022; Yuan et al., 2023; Wang et al., 2024; Jaiswal et al., 2024; Saha et al., 2024; Kaushal et al., 2023; Sharma et al., 2023; Qinsi et al., ) highlight the effectiveness of low-rank decomposition in compressing LLMs. However, despite their flexibility, these methods lag behind semi-structured pruning in performance, often leading to a 2$\times$ increase in perplexity (PPL) at the same densities.
@@ 3.2 Pivoting Factorization
Algorithm 1 Pivoting Factorization
0:  Low-rank matrix $\mathbf{W}^{\prime}\in\mathbb{R}^{m\times n}$ with rank $r$
1:  Use QR (or LU) decomposition with pivoting to find pivot-row indices: $\mathcal{I}\leftarrow\mathrm{QR}_{\text{pivot}}(\mathbf{W}^{\prime})$
2:  Define $\mathcal{I}^{c}=\{1,2,\dots,m\}\setminus\mathcal{I}$, the complement of $\mathcal{I}$, representing non-pivot row indices
3:  Compute pivot-row matrix: $\mathbf{W}_{p}\leftarrow\mathbf{W}^{\prime}[\mathcal{I},:]$
4:  Compute non-pivot-row matrix: $\mathbf{W}_{np}\leftarrow\mathbf{W}^{\prime}[\mathcal{I}^{c},:]$
5:  Compute coefficient matrix $\mathbf{C}$ by solving matrix equation: $\mathbf{C}\leftarrow\mathrm{linsolve}(\mathbf{W}_{np}=\mathbf{C}\mathbf{W}_{p})$
5:  PIFA layer $P$: 1) Pivot-row indices $\mathcal{I}\in\mathbb{R}^{r}$; 2) pivot-row matrix $\mathbf{W}_{p}\in\mathbb{R}^{r\times n}$; 3) coefficient matrix $\mathbf{C}\in\mathbb{R}^{(m-r)\times r}$ for non-pivot rows
@@ 4 Online Error-Accumulation-Minimization Reconstruction
In addition to PIFA, we propose a novel Online Error-Accumulation-Minimization Reconstruction (M) method (illustrated in Figure 2(a)) to reconstruct the low-rank matrix before applying PIFA.
@@ 5 Experiments
We combine Online Error-Accumulation-Minimization Reconstruction with Pivoting Factorization into an end-to-end low-rank compression method, MPIFA (illustrated in Figure 2). MPIFA proceeds as follows: ① First, Online Error-Accumulation-Minimization Reconstruction is applied to obtain and refine the low-rank matrices $\vec{U}_{r}$ and $\vec{V}_{r}^{\mathrm{T}}$; ② Then, PIFA decomposes the singular matrix $\mathbf{W}^{\prime}=\vec{U}_{r}\vec{V}_{r}^{\mathrm{T}}$ into $\mathcal{I},\mathbf{W}_{p},\mathbf{C}\leftarrow\mathrm{PIFA}(\mathbf{W}^{\prime})$, which are stored in a PIFA layer that replaces the original linear layer.
@@
Compared to semi-structured sparsity, MPIFA offers the advantage of allowing non-uniform sparsity for each module. We term the Non-uniform Sparsity version of MPIFA as MPIFANS. The detailed implementation of MPIFANS can be found in Appendix B.2.
@@
We apply MPIFA to pre-trained LLMs: LLaMA2 (7B, 13B, 70B) (Touvron et al., 2023b) and LLaMA3 (8B) (Dubey et al., 2024). Both the calibration data and perplexity (PPL) evaluations are based on the WikiText2 dataset (Merity et al., 2022), with a sequence length of 2048 tokens for all experiments.
@@ 5.1 Main Result
We evaluate MPIFA against state-of-the-art low-rank pruning methods: ASVD (Yuan et al., 2023) and SVD-LLM (Wang et al., 2024). Vanilla SVD is also included for reference. SVD-LLM has two versions, as detailed in their original paper. Following their approach, we select the best-performing version for each density. Results for both versions of SVD-LLM are provided in Table 5.3. MPIFA utilizes 128 calibration samples and sets $\lambda=0.25$. For all models except LLaMA-2-70B, MPIFA reconstructs both $\vec{U}$ and $\vec{V}^{\mathrm{T}}$. For LLaMA-2-70B, MPIFA reconstructs only $\vec{U}$.
@@ 5.3 Ablation Study
[…]
tableAblation: Impact of PIFA and M on perplexity ($\downarrow$) across parameter densities. on WikiText2. W denotes using SVD-LLM’s pruning only; W + U denotes using SVD-LLM’s pruning and full-batch reconstruction; W + M denotes using our Online Error-Accumulation-Minimization Reconstruction, which incorporates SVD-LLM’s pruning as the initial step; W + M + PIFA denotes using Online Error-Accumulation-Minimization Reconstruction followed by PIFA, i.e., MPIFA. Density Model Method 100% 90% 80% 70% 60% 50% 40% LLaMA2-7B W 5.47 7.27 8.38 10.66 16.14 33.27 89.98 W + U 7.60 8.84 11.15 16.11 27.19 54.20 W + M 6.71 7.50 8.86 11.45 16.55 25.26 W + M + PIFA (MPIFA) 5.69 6.16 7.05 8.81 12.77 21.25 LLaMA2-13B W 4.88 5.94 6.66 8.00 10.79 18.38 43.92 W + U 6.45 7.37 9.07 12.52 20.95 42.79 W + M 5.80 6.41 7.42 9.31 13.09 19.93 W + M + PIFA (MPIFA) 5.03 5.39 7.12 7.41 10.30 16.72 LLaMA2-70B W 3.32 4.12 4.58 5.31 6.60 9.09 14.82 W + U 8.23 8.33 8.66 10.02 13.41 22.39 W + M 4.15 4.63 5.31 6.46 8.72 14.11 W + M + PIFA (MPIFA) 3.54 3.96 4.58 5.54 7.40 12.01 LLaMA3-8B W 6.14 9.83 13.62 25.43 76.86 290.3 676.7 W + U 10.63 14.66 23.66 42.60 83.46 163.5 W + M 9.16 11.25 15.27 23.55 36.14 53.85 W + M + PIFA (MPIFA) 6.93 8.31 10.83 16.41 28.90 47.02
@@
Table 5.3 presents an ablation study that evaluates the impact of our Online Error-Accumulation-Minimization Reconstruction (denoted as M) and Pivoting Factorization (PIFA) on perplexity across varying parameter densities. The methods compared include:
@@
Figure 5: Impact of mix ratio. With 0.5 density, MPIFA achieves lowest PPL when the mix ratio $\lambda$ in Equation 7 is around 0.25.
@@
(a) LLaMA2-7B
(b) LLaMA2-13B
Figure 6: Impact of calibration sample size. On MPIFA with 0.5 density, reconstructing both $\mathbf{U}$ and $\vec{V}^{\mathrm{T}}$ is more sensitive to the number of calibration samples than reconstructing only $\mathbf{U}$.
@@
Figures 6a and 6b compare the effects of reconstructing only $\mathbf{U}$, only $\vec{V}^{\mathrm{T}}$, and both $\mathbf{U}$ and $\vec{V}^{\mathrm{T}}$ across different calibration sizes. The results indicate that with sufficient calibration samples, reconstructing both $\mathbf{U}$ and $\vec{V}^{\mathrm{T}}$ achieves lower perplexity than reconstructing only $\mathbf{U}$ or only $\vec{V}^{\mathrm{T}}$.
@@ 6 Conclusion and Discussion
Acknowledgements
This work was supported by the Zhou Yahui Chair Professorship award of Tsinghua University (to CVC), the National High-Level Talent Program of the Ministry of Science and Technology of China (grant number 20241710001, to CVC).
@@ Acknowledgements
Appendix
(a) FP32 Inference Speedup
(b) FP16 Inference Speedup
(c) FP32 Inference Memory
(d) FP16 Inference Memory
Figure 7: Efficiency of PIFA layer under various ranks, with sequence length = 2048, batch size = 32, and dimension = 8192 on FP32 and FP16 on A6000 GPU. At $50\%$ density, PIFA achieves $47.6\%$ memory savings and 1.95×1.95$\times$1.95 × speedup on FP32. These results guarantee that the overhead of both time and memory of the PIFA layer is quite low.
Figure 8: Condition number. Condition numbers of $\vec{V}^{\mathrm{T}}\mathbf{X}\mathbf{X}^{\mathrm{T}}\vec{V}$ (Equation 5) and $\mathbf{X}\mathbf{X}^{\mathrm{T}}$ (Equation 8) for the first layer of LLaMA2-7B, whose inverses are used in reconstructing $\vec{U}$ and $\vec{V}^{\mathrm{T}}$. Larger calibration sizes reduce condition numbers, enhancing numerical stability and lowering perplexity.
Table 3: Efficiency of PIFA layer and semi-sparse layer across different dimensions, compared to dense linear at same dimension on same GPU, with sequence length of 2048 and batch size of 32, using FP16 (FP32 is not supported by 2:4 sparsity in torch.sparse). The highest speedup and lowest memory are indicated in bold. PIFA shows increasing speedup as the dimension grows. For matrix multiplication with weight matrix shape 32768$\times$32768, cuSPARSELt raises CUDA error.
Dimension
GPU Kernel 32768 16384 8192 4096
Speedup A6000 2:4 (cuSPARSELt) Error 0.94$\times$ 0.97$\times$ 1.09$\times$
2:4 (CUTLASS) 0.79$\times$ 0.92$\times$ 1.15$\times$ 1.18$\times$
PIFA 55% 2.10$\times$ 1.88$\times$ 1.70$\times$ 1.43$\times$
A100 2:4 (cuSPARSELt) Error 1.19$\times$ 1.31$\times$ 1.68$\times$
2:4 (CUTLASS) 1.19$\times$ 1.12$\times$ 1.09$\times$ 1.52$\times$
PIFA 55% 1.98$\times$ 1.70$\times$ 1.54$\times$ 1.37$\times$
Memory 2:4 (cuSPARSELt / CUTLASS) 0.564 0.569 0.589 0.651
PIFA 55% 0.552 0.558 0.578 0.645
Table 4: End-to-end efficiency of MPIFANS on LLaMA2 models, on FP16 (FP32 isn’t supported by semi-sparse). The highest throughput and lowest memory are indicated in bold. MPIFANS consistently outperforms semi-sparse in both throughput and memory with 55% density. Enabling KV cache for semi-sparse model will cause SparseSemiStructuredTensorCUSPARSELT only supports a specific set of operations, can’t perform requested op (expand.default)
Model Metrics GPU Use KV Cache Dense 2:4 (cuSPARSELt) 2:4 (CUTLASS) MPIFANS 55%
llama2-7b Throughput (token/s) A6000 No 354.9 306.6 327.5 472.6
Yes 3409 Error Error 4840
A100 No 614.8 636.2 582.3 822.2
Yes 6918 Error Error 7324
Memory (GB) 12.55 7.274 7.290 7.174
llama2-13b Throughput (token/s) A6000 No 190.0 163.2 180.0 268.7
Yes 2156 Error Error 2721
A100 No 345.4 362.4 321.5 473.3
Yes 4217 Error Error 4532
Memory (GB) 24.36 13.90 13.99 13.69
@@ B.3 MPIFANS Fine-tuning
[…]
tableC4 Perplexity ($\downarrow$) at different parameter density (proportion of remaining parameters relative to the original model). The best-performing method is highlighted in bold. Density Model Method 100% 90% 80% 70% 60% 50% 40% LLaMA2-7B SVD 7.29 18931 27154 37208 56751 58451 70567 ASVD 7.98 12.46 201.0 9167 25441 24290 SVD-LLM 13.95 19.89 33.02 61.97 129.8 262.9 MPIFA 8.15 10.20 14.68 25.43 52.01 97.71 LLaMA2-13B SVD 6.74 1994 6301 37250 22783 18196 84680 ASVD 7.15 9.30 23.54 468.5 3537 3703 SVD-LLM 10.93 14.99 24.44 46.65 110.4 267.8 MPIFA 7.27 8.79 21.00 21.33 42.03 80.47 LLaMA2-70B SVD 5.74 10.16 23.28 121.4 1659 7045 12039 ASVD OOM OOM OOM OOM OOM OOM SVD-LLM 7.12 8.71 12.21 21.40 44.10 103.3 MPIFA 6.00 6.76 8.67 13.60 29.04 63.38 LLaMA3-8B SVD 9.47 323597 461991 172968 70896 143573 271176 ASVD 14.43 272.1 8511 18701 108117 9466 SVD-LLM 38.54 98.65 223.5 460.0 784.8 1416 MPIFA 14.76 22.45 44.62 123.0 257.4 429.2
Appendix C Perplexity Evaluation on C4 Dataset
To complement our evaluation on WikiText2, we assess MPIFA on the C4 dataset (Raffel et al., 2020), a widely used benchmark for evaluating LLM perplexity on large-scale, real-world text distributions. For consistency with the main text, we use WikiText2 as the calibration dataset for all methods in this evaluation. The results, summarized in Table B.3, demonstrate that MPIFA consistently achieves the lowest perplexity across most density levels and model sizes, reducing the perplexity gap by 47.6% on LLaMA2-7B, 34.5% on LLaMA2-13B, 55.3% on LLaMA2-70B, and 62.6% on LLaMA3-8B on average across all densities, compared to the best-performing baseline.
Appendix D Zero-Shot Evaluation on SuperGLUE Benchmark
To provide a more comprehensive evaluation of MPIFA’s performance across different compression levels, we conduct zero-shot evaluations at various compression rates on the SuperGLUE benchmark (Wang et al., 2019) using the LLaMA2-7B model.
Evaluations are conducted using the lm-evaluation-harness framework (Gao et al., 2023). All tasks are evaluated using accuracy ($\uparrow$), except for the ReCoRD task, which uses F1 score. The detailed accuracy results are presented in Table 5. The best-performing method at each setting is highlighted in bold. MPIFA consistently achieves the highest mean accuracy across all density levels, outperforming other low-rank methods across a wide range of tasks.
Table 5: Zero-shot evaluations on SuperGLUE datasets at different parameter density on LLaMA2-7B. All tasks are evaluated using accuracy ($\uparrow$), except for the ReCoRD task, which uses F1 score ($\uparrow$). The best-performing method is highlighted in bold.
Density Method BoolQ CB COPA MultiRC ReCoRD RTE WIC WSC Mean
100% Dense 77.7 42.9 87.0 57.0 91.6 63.2 49.7 36.5 63.2
90% SVD 42.6 39.3 67.0 51.0 16.4 55.6 49.8 62.5 48.0
ASVD 55.9 37.5 69.0 47.1 42.5 53.4 49.8 41.3 49.6
SVD-LLM 49.1 41.1 79.0 57.1 87.8 52.7 48.0 48.1 57.8
MPIFA 74.4 64.3 86.0 56.7 91.2 58.5 49.7 36.5 64.7
80% SVD 45.9 57.1 59.0 48.9 12.5 47.3 50.0 58.7 47.4
ASVD 41.6 33.9 58.0 46.8 24.6 55.2 50.0 60.6 46.3
SVD-LLM 44.2 41.1 79.0 55.7 84.7 53.1 50.6 57.7 58.3
MPIFA 69.4 41.1 83.0 55.8 90.3 53.4 48.7 36.5 59.8
70% SVD 40.2 46.4 62.0 43.1 12.4 53.8 48.9 63.5 46.3
ASVD 39.2 46.4 59.0 48.2 14.0 49.1 50.3 63.5 46.2
SVD-LLM 44.4 39.3 82.0 44.6 79.8 53.8 48.9 60.6 56.7
MPIFA 64.8 41.1 80.0 57.2 87.5 53.8 49.1 42.3 59.5
60% SVD 45.5 41.1 61.0 49.4 11.4 51.6 50.0 60.6 46.3
ASVD 48.9 37.5 59.0 46.6 15.2 50.2 50.0 40.4 43.5
SVD-LLM 38.5 39.3 71.0 44.9 66.4 53.4 50.2 59.6 52.9
MPIFA 56.6 35.7 79.0 44.5 82.8 57.8 52.0 63.5 59.0
50% SVD 38.6 44.6 63.0 43.0 10.1 52.7 50.2 63.5 45.7
ASVD 42.2 39.3 63.0 45.7 14.4 52.7 50.0 63.5 46.3
SVD-LLM 37.9 41.1 66.0 42.8 50.5 52.7 50.0 63.5 50.5
MPIFA 38.0 35.7 70.0 42.8 71.1 52.4 50.0 63.5 52.9
40% SVD 49.5 42.9 61.0 48.1 11.3 52.0 50.0 49.0 45.5
ASVD 42.6 42.9 59.0 47.7 14.5 53.1 50.0 64.4 46.8
SVD-LLM 37.8 41.1 67.0 42.8 37.0 52.7 50.0 63.5 49.0
MPIFA 37.8 37.5 68.0 42.8 54.7 53.8 50.0 63.5 51.0
Appendix E Comparison with Structured Pruning Baseline
To provide a fair comparison with structured pruning methods, we include additional evaluations of LLM-Pruner (Ma et al., 2023) as a baseline. All experiments are conducted on the WikiText2 dataset using the LLaMA2-7B model.
Perplexity Comparison.
Table 6 shows the perplexity across various parameter densities. On average, MPIFA reduces the perplexity gap by 87.2% compared to LLM-Pruner.
Table 6: LLM-Pruner vs. MPIFA: Perplexity Comparison on WikiText2 using LLaMA2-7B at different densities. MPIFA consistently achieves lower perplexity across all densities.
Method 100% 90% 80% 70% 60% 50% 40%
LLM-Pruner 5.47 6.58 8.81 13.70 40.49 126.0 1042
MPIFA 5.69 6.16 7.05 8.81 12.77 21.25
Inference Speedup and Memory Efficiency.
Table 7 and Table 8 compare the inference speedup and memory usage (relative to dense linear layer) for PIFA and LLM-Pruner layers, across different hidden dimensions on an A6000 GPU.
Table 7: Inference speedup (× over dense) for PIFA and LLM-Pruner layers.
Method (density) d=16384 d=8192 d=4096
PIFA (55%) 1.88× 1.70× 1.43×
LLM-Pruner (55%) 1.81× 1.77× 1.67×
LLM-Pruner (70%) 1.42× 1.41× 1.35×
Table 8: Memory usage (× over dense) for PIFA and LLM-Pruner layers.
Method (density) d=16384 d=8192 d=4096
PIFA (55%) 0.56× 0.58× 0.64×
LLM-Pruner (55%) 0.56× 0.58× 0.65×
LLM-Pruner (70%) 0.70× 0.72× 0.75×
At the same density (55%), PIFA achieves similar speedup and memory efficiency as LLM-Pruner. When comparing MPIFA at 55% density to LLM-Pruner at 70% density, MPIFA consistently achieves lower perplexity, faster inference, and reduced memory usage.
Appendix F Compression Time and Memory Usage Comparison
We provide a detailed comparison of the compression time and peak GPU memory usage during compression across different methods. These metrics reflect the practical resource requirements of each approach.
Compression Time.
Table 9 reports the time required to perform compression for each method, measured on a single A6000 GPU. On an A100 GPU, the compression time is approximately halved. ASVD requires significantly longer compression time, up to 10–20 hours, due to the need for sensitivity-based truncation rank searching. The M method involves only reconstruction.
Table 9: Compression time for different methods on a single A6000 GPU.
Model ASVD SVD-LLM PIFA M
LLaMA2-7B 10h 30 min 15 min 30 min
LLaMA2-13B 20h 1h 30 min 1h
Peak Memory Usage During Compression.
Table 10 reports the peak GPU memory usage during compression. Method M has a relatively low memory footprint, as it processes one layer and one sample at a time.
Table 10: Peak GPU memory usage (GB) during compression for different methods.
Model ASVD SVD-LLM PIFA M
LLaMA2-7B 15G 20G 0.5G 6G
LLaMA2-13B 30G 25G 1G 10G
For the M method, we report only the reconstruction time, excluding the time required for the initial low-rank pruning step, which could be SVD or SVD-LLM. The low memory footprint of M is primarily attributed to:
  • Online calibration: Only the current sample is loaded into GPU memory, while the rest of the samples remain in CPU memory until needed.
  • Layer-wise loading: Only the current pruning layer is loaded to the GPU at any time, with all other layers remaining on CPU.
Appendix G Enhancing Other Low-Rank Pruning Methods with PIFA and M
To demonstrate the generality of our proposed methods, we evaluate how PIFA and M can be applied on top of existing low-rank pruning techniques beyond SVD-LLM, including the pruning strategies proposed in the ESPACE paper (Sakr & Khailany, ).
For a pruning-only comparison, we reproduce the pruning step of ESPACE and compare its performance when combined with PIFA and M. ESPACE introduces six variants: MSE (Eq. 6), MSE-NORM (Eq. 7), GO-MSE (Eq. 8), GO-MSE-NORM (Eq. 8), NL-MSE (Eq. 9), and NL-MSE-NORM (Eq. 9). We exclude the NL-MSE variants as they rely on backpropagation, which is infeasible on memory-constrained GPUs.
Table 11 reports the perplexity (PPL) on WikiText2 at 50% density using the LLaMA2-7B model, showing how PIFA and M further improve various low-rank pruning baselines.
Table 11: Perplexity (PPL) on WikiText2 at 50% density using LLaMA2-7B.
Pruning Method (X) X X + PIFA X + M X + MPIFA
SVD-LLM (W) 33.27 19.64 16.55 12.77
ESPACE (MSE) 280.19 144.32 20.99 16.84
ESPACE (MSE-NORM) 172.30 113.82 21.73 17.42
ESPACE (GO-MSE) 41.75 24.17 17.47 13.55
ESPACE (GO-MSE-NORM) 37.45 23.19 17.47 13.63
Here, SVD-LLM (W) refers to the standalone pruning output from SVD-LLM without additional reconstruction.
PIFA and M consistently improve the performance of different low-rank pruning methods, including ESPACE, demonstrating that they are general-purpose techniques applicable across various low-rank strategies. These results suggest that PIFA and M are not specific to SVD-LLM but can serve as plug-in modules for other pruning methods to improve accuracy while maintaining efficiency.